OpenStudy (lina17):

solve for x : -3|2x-6|= -12

pooja195 (pooja195):

-3|2x-6|= -12 divide both sides by 3 now we have this |2x−6|=4 an absolute value has 2 solutions one positive and one negative 2x-6=4 2x-6=-4 solve both for x

OpenStudy (lina17):

? im kinda lost

pooja195 (pooja195):

on which part?

OpenStudy (lina17):

the sovle both for x what em i suppose to do

pooja195 (pooja195):

2x-6=4 simply isolate it start by adding 6 to both sides -6+6=? 4+6=?

OpenStudy (lina17):

so then you would get 2x=2 ?

OpenStudy (lina17):

or theres no solution

OpenStudy (lina17):

?

pooja195 (pooja195):

Not quite \[\huge~\rm~ 2x−6+6=4+6\] \[\huge~\rm~ 2x=10\] divide both sides by 2 x= ?

OpenStudy (lina17):

so x =5

pooja195 (pooja195):

Good now for the 2nd solution you try it! :) 2x-6=-4

OpenStudy (lina17):

isnt that the one we just did ?

pooja195 (pooja195):

it has a negative 4 keep in mind that an absolute value has 2 solutions

OpenStudy (lina17):

oh okay let me try

OpenStudy (lina17):

so it would just be x

OpenStudy (lina17):

x=0

OpenStudy (lina17):

or i mean x=1

pooja195 (pooja195):

correct :)

pooja195 (pooja195):

x= 1

pooja195 (pooja195):

So what would our 2 answers be?

OpenStudy (lina17):

thanks can you help me on one more problem

pooja195 (pooja195):

sure :)

OpenStudy (lina17):

its |x-2|+10=12

pooja195 (pooja195):

ok any ideas on what to do first?

OpenStudy (lina17):

do i subtract 10 on each side

OpenStudy (lina17):

?

pooja195 (pooja195):

*yes

OpenStudy (lina17):

because i have to isolate the x

OpenStudy (lina17):

ok the answer i got was x= -4 and x=0

pooja195 (pooja195):

correct set up the two problems x-2=2 x-2=-2 Perfect :D welll done!

OpenStudy (lina17):

thank you

pooja195 (pooja195):

yw ^_^